can someone help me understand how to solve this type of problems thanks...
Problem #1
Using the union rule to answer the following questions.
suppose n(A (A LETTER "u" UPSIDE DOWN) b) = 5
n(A U B) = 35
n (A) = 13
what is n(B)?

Problem #2
Show that the stamtements is true by drawing venn diagrams and shading the regions representing the sets on each side of the equals sign.
(A (UPSIDE DOWN U) B)' = A' U B'

Answer by stanbon(57347)

Problem #1
Using the union rule to answer the following questions.
suppose n(A (A LETTER "u" UPSIDE DOWN) b) = 5
n(A U B) = 35
n (A) = 13
what is n(B)?
Ans:
n(A union B) = n(A)+n(B)-n(A intersect B)
35 = 13 + n(B) - 5
n(B) = 35-8
n(B) = 27
======================

Problem #2
Show that the statements is true by drawing venn diagrams and shading the regions representing the sets on each side of the equals sign.
(A (UPSIDE DOWN U) B)' = A' U B'
Draw two intersecting circles in a rectangle.
Draw it again.
Shade (A intersect B)' in the 1st
Shade (A' union B') in the 2nd
------
Do the two shaded areas look the same?
They should. This is one piece of DeMoivre's Rule.
============
Cheers,
Stan H.